which of the following could be the equation of a line thats parallel to 18x-72y+90=0 and passing through the mid point b/w (-14,10) and (10,-2) ?
a)4x-16y+48=0
b)x-4y+18=0
c)-3x+12y+36=0
d)-7x+28y-122=0

Answer by lwsshak3(11628)

which of the following could be the equation of a line thats parallel to 18x-72y+90=0 and passing through the mid point b/w (-14,10) and (10,-2) ?
**
coordinates of midpoint:
(x1+x2/2,(y1+y2)/2
=(10-14)/2,(-2+10)/2
=(-2,4)
..
18x-72y+90=0
x-4y+5=0
4y=x+5
y=x/4+5/4
slope,m=1/4
.
equation:
y=x/4+b
solve for b using coordinates of the midpoint (-2,4)
4=-2/4+b
b=9/2
equation of line parallel to given line and passing thru midpoint.
y=x/4+9/2
4y=x+18
x-4y+18=0 (ans b)